An object is launched from ground level directly upward at a rate of 144
feet/sec. The equation for object’s height is
2
y x x = − + 16 144
. What values of x is the
object at or above a height of 288 feet?

Answer by Alan3354(69443)

An object is launched from ground level directly upward at a rate of 144
feet/sec. The equation for object’s height is
2
y x x = − + 16 144
. What values of x is the
object at or above a height of 288 feet?
-----------------
h(t) = -16t^2 + 144t is a usable equation.
-16t^2 + 144t = 288
t^2 - 9t + 18 = 0
(t-3)*(t-6) = 0
t = 3 and t = 6 seconds